Hawkes Learring nbox (505) bmit Assignment Pause Ch3Sect Question 11of 18 Step 1 of 1 Gilven the following weights and data vallues, calculate the mean. Round your answer to three decimal places Weights1.79 2.45 5.60 4.39 2.41 381 5.37 7.05 Data Values 2 7 63 9 6 Answer How to Enter) 1 Point Keypad
Answer #1
c weights wi*xi
2 1.79 3.2041
2 2.45 6.0025
7 5.6 31.36
6 4.39 19.2721
3 2.41 5.8081
9 3.81 14.5161
1 5.37 28.8369
6 7.05 49.7025
32.87 158.7023
mean 4.828180712
c weights wi*xi
2 1.79 =B2*B2
2 2.45 =B3*B3
7 5.6 =B4*B4
6 4.39 =B5*B5
3 2.41 =B6*B6
9 3.81 =B7*B7
1 5.37 =B8*B8
6 7.05 =B9*B9
=SUM(B2:B9) =SUM(C2:C9)
mean =C11/B11

mean =4.828180712

= 4.828

Please rate if you found this response useful

Please comment for any further clarification/discussion